OpenStudy (eli_moses):

(please help!! fan and medal!! i don't know where to start!!) H(x)=(x^2)+1 and K(x)=-(x^2)+4. If K(H)=0, what are the roots/solutions? a. ±i√3 b. ±i c. ±2 d. ±1 e. ±1, ±i√3

OpenStudy (anonymous):

\[ K(H) = -(H^2)+4 = -((x^2+1)^2)+4 \]Simplify.

OpenStudy (anonymous):

I don't even know!

OpenStudy (anonymous):

Then you can find roots.

OpenStudy (jdoe0001):

\(\bf K(H) = -(H^2)+4 = -[(x^2+1)^2]+4\to \begin{cases} 4-[(x^2+1)^2] \\ \quad \\ \bf {\color{brown}{ a^2-b^2 = (a-b)(a+b)}} \end{cases}\)

OpenStudy (phi):

As people have pointed out K(H) means wherever you see an x in K, replace it with x^2 +1 and then set the expression equal to zero. you get \[ -(x^2+1)^2 + 4 =0 \] add -(x^2+1)^2 to both sides to get \[ (x^2+1)^2 =4 \] the next step is to take the square root of both sides. what do you get ?

OpenStudy (eli_moses):

\[x ^{2}+1=2\]

OpenStudy (phi):

almost, you get \( x^2 + 1= \pm 2\) now you have to solve two equations \( x^2 + 1=2 \\x^2 + 1=-2 \)

OpenStudy (phi):

can you solve the first equation (you get two answers for x)

OpenStudy (eli_moses):

1 and -1?

OpenStudy (phi):

yes. now do the second equation

OpenStudy (phi):

you should get \[ x^2 = -3 \\ x= \pm \sqrt{-3} \] that is normally written using \( i = \sqrt{-1} \), as \[ x= \pm i \sqrt{3} \]

OpenStudy (eli_moses):

i was about to ask that

OpenStudy (phi):

we use the idea that \[ \sqrt{a \cdot b} = \sqrt{a} \sqrt{b} \] and the definition \( i = \sqrt{-1} \), to write \[ \sqrt{-3} = \sqrt{-1 \cdot 3} = \sqrt{-1} \sqrt{3} \\ = i \sqrt{3} \]

OpenStudy (eli_moses):

so the answer would be a?

OpenStudy (phi):

you have 4 solutions (did you forget about +1 and -1 ) ?

OpenStudy (eli_moses):

oh, right.

OpenStudy (eli_moses):

so the answer is choice e, right?

OpenStudy (phi):

yes, choice E

OpenStudy (eli_moses):

thanks phi

OpenStudy (phi):

you can check all 4 solutions. for example, when x=1 we first do H(1)=(1^2)+1 = 1+1= 2 now do K(2) = -(2^2)+4 = -4 + 4 = 0 (it works) here is x= -i sqr(3) \[ H(-i \sqrt{3})=(-i \sqrt{3})^2+1 = -i \cdot -i \cdot 3+1= i^2\cdot 3 +1= -3 +1= -2\] and K(-2)= -(-2)^2 + 4 = - 4 + 4 = 0 (also works)

OpenStudy (eli_moses):

thanks